OpenStudy (anonymous):

The average of 4 silbings is 12.5 years if 3 of the silbings are 12 years old triplets how old is the fouth silbing????

OpenStudy (anonymous):

let the age of each of the siblings be a,b,c,d respectively. then the average age is (a+b+c+d)/4 = 12.5 now 3 are 12, so plug these values in, (12+12+12+d)/4=12.5 36+d=50 d=14

OpenStudy (anonymous):

\[\frac{12+12+12+x}{4}=12.5\]

OpenStudy (anonymous):

bee is right thx